# MC-MKE: A Fine-Grained Multimodal Knowledge Editing Benchmark Emphasizing Modality Consistency

### v1

@@ 5 Experiments
Mini-GPTv2 FT(Vision) 86.86 3.12 86.33 52.87 6.01
FT(LLM) 95.82 2.40 94.30 94.65 87.35
MEND(Vision) 6.61 50.06 6.06 7.61 2.74
MEND(LLM) 63.13 46.55 59.79 38.77 5.09
IKE 17.51 / 38.57 / 11.09
Table 5: Experimental results on IRO_edit data for three editing methods editing two different model components on two MLLMs.
5.1 Multimodal Large Language Models
InstructBLIP InstructBLIP is a multimodal large language model that consists of three modules. Its multimodal alignment module consists of a Qformer structure and a linear layer network to connect its vision and large language model module. We use InstructBLIP equipped with Vicuna-7B Chiang et al. (2023).
MiniGPT-v2 MiniGPT-v2 utilizes a linear projection layer as an alignment module to map visual features to LLM feature space. Compared with InstructBLIP, MiniGPT-v2 has a smaller alignment module but still more input visual features. We use MiniGPT-v2 equipped with Llama-2-Chat-7B Touvron et al. (2023).
5.2 MMEdit Methods
There have been many language knowledge editing methods, while multi-modality knowledge editing methods have not been fully explored. Therefore, we select the following representative editing methods in multimodal knowledge editing according to the editing requirements of the task and the applicability of the methods.
Finetune
Finetune is one of the most widely used and apparent method to improving or modifying the abilities of pre-trained models and is also generally used as a baseline for knowledge editing. Since one can select the model component to finetune, it is natural to explore the differences between finetuning different model components. We focus on finetuning two parts: the alignment module and the LLM component of an MLLM. For LLM component, we only finetune the last layer of the LLM.
MEND
Model Editor Networks with Gradient Decomposition (MEND) (Mitchell et al., 2022) is a editor network mapping a single desired input-output knowledge pair to the corresponding parameter update of the original model. Specifically, the input-output knowledge pair provides a standard fine-tuning gradient as a starting point for editing updates. Then MEND directly transforms the gradient to a better parameter update ensuring both generality and locality.
IKE
In-Context Knowledge Editing (IKE) (Zheng et al., 2023) enables knowledge editing by incorporating demonstration examples within the input data to update and acquire new factual knowledge without the requirement of further training. Considering the instruction-following ability of the MLLM and the limitation on the number of input images, we choose to implement the zero-shot model of IKE. More experimental details can be seen in appendix B.
@@ 5.3 Results & Analysis
Pros and Cons of Different Editing Methods We observe that across all editing formats and model modules, no existing editing method perfectly meets our editing requirements.
The Finetune method is characterized by high Reliability, demonstrating good Generality and Consistency when editing the LLM part in SRO_edit and IRO_edit. However, its Locality is very low, meaning it has a significant impact on unrelated knowledge.
MEND, although it also modifies the model’s parameters, uses a meta-learning approach to control the model’s changes to other unrelated knowledge. As shown from the results, MEND’s Reliability is much lower than that of Finetune, but its Locality is extremely higher.
As for IKE, since many MLLMs do not support in-context learning for image inputs, we do not test it for Locality and M-Generality in IE_edit and IRO_edit. Since it relies on in-context learning, this method is inherently sensitive to prompts. Different types of editing formats correspond to different prompts, and different models have varying degrees of sensitivity to prompts, resulting in significant fluctuations in all of IKE’s metrics. Overall, IKE performs better on InstructBLIP, achieving the highest Locality in SRO_edit. It also shows high Consistency on InstructBLIP, achieving the highest Consistency metric in IE_edit, indicating that the model can infer the edited multimodal knowledge $(i,\tilde{r},\tilde{o})$ based on context and the image.
@@
Consistency On Different Editing Formats In SRO_edit and IRO_edit, the output of their corresponding Consistency test data matches the required edited output, with only the input information being different. Therefore, if the model can correctly understand the different formats of questions, it can accurately answer the Consistency questions. In these two editing formats, Finetune and IKE achieve high Consistency on InstructBLIP. However, on MiniGPT-v2, these methods only maintain some degree of Consistency. In these two editing formats, high Consistency without high Locality may come from overfitting. Thus, to accurately assess the Consistency property, we need to analyze the IE_edit format. On InstructBLIP, the Finetune method maintains high Consistency with high Reliability, indicating that the Finetune method is not solely overfitting since $\tilde{e}\neq\tilde{o}$. Only when a method achieves high Consistency across all three editing formats can its Consistency property be considered trustworthy. Considering the results across the two models, IKE shows best Consistency, managing to maintain high Consistency while ensuring a certain degree of Locality.
Editing Different Components Cheng et al. (2024) mentioned the visual module is harder to edit compared to the text module. In SRO_edit and IRO_edit, apart from Locality, the effectiveness of editing the visual module is much lower than that of editing the LLM module. For Locality, editing the alignment module using MEND has a smaller impact, possibly because MEND’s parameter fitting is not as strong, and the editing knowledge in SRO_edit and IRO_edit is the textual knowledge $(s,r,o)$ triplet. However, in IE_edit, although editing the LLM module generally yields higher Reliability and slightly higher Generality, in the case of InstructBLIP, editing the LLM module has a lower Consistency. This indicates that editing the LLM module often leads to overfitting, where the model outputs the edited knowledge regardless of the input information. In contrast, editing the Vision module, although resulting in lower other metrics, still maintains high Consistency. This shows that in IE_edit, editing the visual module can still ensure the Consistency of the corresponding knowledge.
@@ 6 Conclusion
The main limitations of our work is related to limited knowledge editing methods and multimodal large language models. We only provide results on the two latest MLLMs, including InstructBLIP and MiniGPT-v2, leaving many others behind. As we study the latest MLLMs on knowledge editing methods which have not been discussed in prior work, we only analyze three knowledge editing methods, Finetune, MEND and IKE.
@@
CMKoBe is a synthetic dataset constructed by randomly modifying the factual knowledge triplets, rather than being crafted by humans. The data samples could accidentally involve context which is toxic or offensive in nature. ChatGPT is used for data annotation and assisting writing.
@@ Appendix A Pre-experiments
SRO_edit focuses on editing a textual knowledge triplets $(s,r,o)$, inherently requiring no additional visual inputs. But to align with the standard input format of MLLMs, we input a black image as the visual placeholder. In this section, we present an preliminary experiment to explore different choices of the input visual images including black images, white images and random noise. The results of InstructBLIP with these three types of images on SRO_edit are 95.11, 96.53 and 94.70 respectively. It is shown that these uninformative images barely have influence on the results.
@@ Appendix B Experiment Details
We list the hyper-parameters used for finetuning in Table 6. MiniGPT-v2 and InstructBLIP share the same hyper-parameters.
@@
Training process of MEND requires additional training data specific to the underlying model. Following Mitchell et al. (2022), we construct an edit dataset and a locality dataset for both InstructBLIP and MiniGPT-v2. We leverage the data filtered in Section 4.1 as the edit dataset, sharing identical distribution with MC-MKE. Since both InstructBLIP and MiniGPT-v2 leverage MS COCOLin et al. (2015) for pretraining, we include it as the locality dataset. We search for three important hyper-parameters $c_{loc}$, $c_{edit}$ and learning rate on each experimental setting for ten times. We found that MEND is very sensitive to hyperparameters, especially when the target module is small (e.g. the MEND(Vision) setting in our main experiment).
@@ 5 Rephrased Texts:
input [Uncaptioned image] Question: The country in the picture is
$\tilde{e}$: Lithuania
$p_{r}$ [Uncaptioned image] Question: The country in the picture is
$t_{r}$: Lithuania
Alias: Lietuva, Lietuvos Respublika, …
$p_{c}$ [Uncaptioned image] Question: The capital of the country in the picture is
$t_{c}$: Vilnius
Alias: Vilnia, Vilna, Wilno, …
@@
Question: The country in the picture is
$t^{M}_{g}$: Lithuania
Alias: Lietuva, Lietuvos Respublika, …
@@
Question: Can you tell me which country is depicted in the image?
$t^{T}_{g}$: Lithuania
Alias: Lietuva, Lietuvos Respublika, …
@@
input / Question: The capital of United Kingdom is
$\tilde{o}$: Rupnagar
$p_{r}$ / Question: The capital of United Kingdom is
$t_{r}$: Rupnagar
Alias: Rupar, Ropar
$p_{c}$ [Uncaptioned image] Question: The capital of the country in the picture is
$t_{c}$: Rupnagar
Alias: Rupar, Ropar
@@
$p^{T}_{g}$ / Question: What is the capital of United Kingdom?
$t^{T}_{g}$: Rupnagar
Alias: Rupar, Ropar
@@
input [Uncaptioned image] Question: As a result of World War III, the country in the picture moves its capital. The capital of the country in the picture is
$\tilde{o}$: Rupnagar
$p_{r}$ [Uncaptioned image] Question: The capital of the country in the picture is
$t_{r}$: Rupnagar
Alias: Rupar, Ropar
$p_{c}$ / Question: The capital of United Kingdom is
$t_{c}$: Rupnagar
Alias: Rupar, Ropar
@@
Question: Which TV channel is shown in the picture?
$t_{l}$: English
Alias: English language, …
@@
Question: The capital of the country in the picture is
$t^{M}_{g}$: Rupnagar
Alias: Rupar, Ropar
@@
Question: Can you tell me the capital of the country shown in the picture?
$t^{T}_{g}$: Rupnagar
Alias: Rupar, Ropar

### v2

@@ 5 Experiments
5.1 MMEdit Methods
There have been many language knowledge editing methods, while multi-modality knowledge editing methods have not been fully explored. Therefore, we select the following representative editing methods including Finetuning, MEND (Mitchell et al., 2022a), IKE (Zheng et al., 2023) and SERAC (Mitchell et al., 2022b) in single-modal knowledge editing following previous setting(Cheng et al., 2024). More information and implementation details of these editing methods can be seen in Appendix B.
@@ 5.1 MMEdit Methods
SERAC 91.65 99.06 26.01 91.47 86.46
MiniGPT-v2 FT(Vision) 98.98 73.71 98.78 93.32 24.13
FT(LLM) 88.49 2.04 86.99 87.25 84.32
MEND(Vision) 6.21 76.00 5.45 4.52 2.13
MEND(LLM) 34.21 67.31 43.91 25.49 6.72
IKE 62.73 / 62.48 / 21.49
SERAC 88.49 97.25 26.92 87.25 84.32
Table 5: Experimental results on IRO_edit data for four editing methods editing two different model components on two MLLMs. The highest value is highlighted in bold.
@@ 5.2 Results & Analysis
Consistency On Different Editing Formats
@@
In SRO_edit and IRO_edit, the output of their corresponding Consistency output matches the required edited output, with only the input information being different. In these two editing formats, high Consistency without high Locality may come from overfitting. Thus, to accurately assess the Consistency property, we need to analyze the IE_edit format as well, where the Consistency output $\tilde{o}$ is different from edited output $\tilde{e}$.
According to Table 2, the FT(Vision) maintains high Consistency with InstructBLIP, indicating that the FT(Vision) is not solely overfitting to obtain Consistency in IE_edit dataset. Only when a method achieves high Consistency across all three editing formats can its Consistency property be considered trustworthy. Overall, IKE shows good Consistency while maintaining a certain degree of Locality. However, even IKE shows unsatisfactory consistency performance on the IE_edit dataset, and its performance on MiniGPT-v2 on the other two datasets is even worse. This indicates that the current tested methods may ignore Consistency during development, resulting in their inability to maintain high consistency across all datasets.
Pros and Cons of Different Editing Methods
The Finetuning method is characterized by high Reliability. It also demonstrates good Generality and Consistency when editing the LLM part in SRO_edit and IRO_edit. However, its Locality is not satisfying, suggesting it has a significant impact on unrelated knowledge.
MEND employs a meta-learning approach to adjust the model’s parameters while minimizing effects on unrelated knowledge. As shown from the results, MEND bears a much lower Reliability but higher Locality than Finetuning. The meta-learning approach, while preventing the modification of knowledge irrelevant to the model, also reduces the accuracy of editing relevant knowledge. Moreover, the performance of meta-learning is highly unstable across different types of data and models, according to results in Table 4 and Table 5.
As for IKE, since many MLLMs do not support in-context learning for image inputs, we do not test it for Locality and M-Generality in IE_edit and IRO_edit. IKE achieving good Locality in SRO_edit. However, IKE relies on in-context learning, which is inherently sensitive to prompts. Different types of editing formats require different prompts, and different models have varied sensitivity degrees to prompts, resulting in significant fluctuations in all metrics of IKE. Overall, IKE performs better on InstructBLIP. In IE_edit, it also shows high Consistency on InstructBLIP, achieving the highest Consistency. indicating that the model can infer the edited multimodal knowledge $(i,\tilde{r},\tilde{o})$ based on context and the image.
SERAC applies a classifier to choose whether to use the original model or the counterfactual model. SERAC achieves high levels of Reliability and Locality, as well as Generality in most cases. However, its performance relies heavily on the classifier performance, whether the classifier can correctly identify the appropriate model for the given input. Although SERAC obtains good Consistency with good Locality in SRO_edit and IRO_edit datasets, its Consistency in IE_edit is still low. While the classifier can sometimes effectively distinguish between inputs related to edited knowledge and those that are not, it still cannot directly improve Consistency. Even if the classifier identifies the need to use the counterfactual model to answer questions in the Consistency test, the ability to respond to the Consistency test still depends on the counterfactual model itself.
We observe the results across all editing formats and models, no existing editing method perfectly meets all editing requirements. Most methods have problems with consistency, according to Table 2, 4 and 5.
Editing Different Components
Cheng et al. (2024) mentioned the visual module is harder to edit compared to the text module. Based on our experimental results, this point holds true in some cases, but in other instances, editing the visual part may yield better results in certain aspects. For MEND, meta-learning requires predicting network changes corresponding to the knowledge edits, and editing the visual module to output the edited knowledge is more challenging. As a result, in most cases, using MEND(Vision) tends to result in lower reliability.
While the MEND approach does help prevent the modification of irrelevant knowledge to some extent, editing the LLM module with MEND still often achieves lower locality as shown in Table 4 and 5. Furthermore, in the IE_edit dataset, MEND achieves higher consistency when editing the visual module. Across the three datasets, FT(Vision) often achieves reliability similar to FT(LLM). On MiniGPT-v2, FT(Vision) results in higher locality. When editing MiniGPT-v2, although FT(LLM) tends to yield higher consistency on the SRO_edit and IRO_edit datasets, it comes with low locality. The higher consistency brought by FT(LLM) in SRO_edit and IRO_edit datasets may be due to overfitting to the edited output. This is evidenced by the IE_edit dataset, where the consistency output is different from the edited output. In IE_edit dataset, FT(LLM) yields worse consistency than FT(Vision). What’s more, in IE_edit dataset, FT(Vision) achieves the highest Consistency among all parameter-based methods, which may indicate that editing the Vision part may be better when editing visual knowledge.
@@ 6 Conclusion
The main limitations of our work are related to limited knowledge editing methods and multimodal large language models. We only provide results on MLLMs with 7B checkpoint. We were unable to test larger checkpoints, due to resource constraints.As we study the latest MLLMs on four knowledge editing methods which have not been discussed in prior work, we need to implement them from scratch. We end up implement four knowledge editing methods, Finetuning, MEND, IKE and SERAC.
@@
MC-MKE: is a synthetic dataset constructed by randomly modifying the factual knowledge triplets, rather than being crafted by humans. The data samples could accidentally involve context which is toxic or offensive in nature. ChatGPT is used for data annotation and assisting writing.
@@ Appendix A Pre-experiments
SRO_edit focuses on editing a textual knowledge triplets $(s,r,o)$, inherently requiring no additional visual inputs. But to align with the standard input format of MLLMs, we input a black image as the visual placeholder. In this section, we present a preliminary experiment to explore different choices of the input visual images including black images, white images and random noise. The accuracy of InstructBLIP with these three types of images on SRO_edit are 95.11, 96.53 and 94.70 respectively. It is shown that these uninformative images barely have influence on the results.
@@ Appendix B Experiment Details
Finetuning is one of the most widely used and apparent methods for improving or modifying the abilities of pre-trained models and is also generally used as a baseline for knowledge editing. Since one can select the model component to finetune, it is natural to explore the differences between finetuning different model components. We focus on finetuning two parts: the alignment module and the LLM component of an MLLM. For the LLM component, we only finetune the last layer. We list the hyper-parameters used for finetuning in Table 6. MiniGPT-v2 and InstructBLIP share the same hyper-parameters.
@@
Model Editor Networks with Gradient Decomposition (MEND) (Mitchell et al., 2022a) is an editor network mapping a single desired input-output knowledge pair to the corresponding parameter update of the original model. Specifically, the input-output knowledge pair provides a standard fine-tuning gradient as a starting point for editing updates. Then MEND directly transforms the gradient to a better parameter update ensuring both generality and locality. Training process of MEND requires additional training data specific to the underlying model. Following Mitchell et al. (2022a), we construct an edit dataset and a locality dataset for both InstructBLIP and MiniGPT-v2. We leverage the data filtered in Section 4.1 as the edit dataset, sharing identical distribution with MC-MKE. Since both InstructBLIP and MiniGPT-v2 leverage MS COCOLin et al. (2015) for pretraining, we include it as the locality training dataset. We search for three important hyper-parameters $c_{loc}$, $c_{edit}$ and learning rate on each experimental setting for ten times. We found that MEND is very sensitive to hyperparameters, especially when the target module is small (e.g. the MEND(Vision) setting in our main experiment).
IKE Details
In-Context Knowledge Editing (IKE)(Zheng et al., 2023) enables knowledge editing by incorporating demonstration examples within the input data to update and acquire new factual knowledge without the requirement of further training. Considering the limitation on the number of input images, we choose to implement the zero-shot version of IKE.
SERAC Details
SERAC(Mitchell et al., 2022b) proposes a memory-based editing approach. The approach consists of a classifier and a counterfactual model. The classifier chooses whether to use the counterfactual model or not based on the relation between the given input and edit memory.
Since our tasks are multimodal, we use a neural network trained on the training set as the classifier. The neural network consists of a CLIP feature extraction layer and an MLP classification layer. We set the learning rate of the classification layer to 0.0005. Since consistency requires the model to have reasoning abilities, we opted to continue using the large model as the counterfactual model. Specifically, we employ a large model with its LLM part fine-tuned on the edited knowledge as the counterfactual model.
MLLMs Details
InstructBLIP is a multimodal large language model that consists of three modules. Its multimodal alignment module consists of a Qformer structure and a linear layer network to connect its vision and large language model module. We use InstructBLIP equipped with Vicuna-7B Chiang et al. (2023).
MiniGPT-v2 utilizes a linear projection layer as an alignment module to map visual features to LLM feature space. Compared with InstructBLIP, MiniGPT-v2 has a smaller alignment module but still more input visual features. We use MiniGPT-v2 equipped with Llama-2-Chat-7B Touvron et al. (2023).
@@ Appendix C Data Details
Data Selection Details
We filter the data using a completely black image paired with questions in MQuAKE dataset. We selected data that our MLLMs could correctly answer. This step ensures that all the edited knowledge is originally known by the model to make sure we are “editing" instead of “learning". The filtered dataset is referred to as $D_{filter_{1}}$.
Finally, we replaced the subject in the questions with “the {category} in the picture”, seen in Appendix D. If the combined question can be correctly answered by all models, the data is then retained. This step ensures the original multimodal knowledge consistency. The final retained multimodal knowledge constitutes our knowledge editing source dataset $D_{orig}$.
@@
Entity Category Generation Evaluation
We employ ChatGPT to generate the category of a given entity. To verify the quality of categories generated by ChatGPT, we randomly sampled 200 items and invited two annotators to independently verify whether the entities mentioned in these items matched their respective categories. The average agreement between the annotators was 98%, with a consistency rate of 97%, indicating that the generated entity categories are highly reliable. An example of a generated entity category is: "google" : "company".
Training set Construction
Except for not undergoing the original problem filtering, the construction of the train data is similar to that of the test set. We utilize some of the filtered data to construct training set. For the filtered data which are not in $D_{orig}$, we directly apply the question in MQuAKE dataset as text generality textual input, and use the images from google as image generality visual input.
Rephrase Generation Evaluation
We employ ChatGPT to generate Generality data. To verify the quality of rephrases generated by ChatGPT, we randomly sampled 100 items each associated with 4 paraphrased sentences and asked two annotators to independently assess the quality of each paraphrased sentence, marking them as 0 for bad quality and 1 for good quality. The average scores for the 400 paraphrase results were 0.9675, respectively, with an agreement of 98%, demonstrating that the quality of our paraphrases is sufficiently reliable. An example of paraphrased sentence is: Origin : "Who performed Folsom Prison Blues?" Rephrase : "Who was the performer of Folsom Prison Blues?"
@@ 5 Rephrased Texts:
input [Uncaptioned image] Question: The book in the picture is
$\tilde{e}$: The Pilgrim’s Progress
$p_{r}$ [Uncaptioned image] Question: The book in the picture is
$t_{r}$: The Pilgrim’s Progress
Alias: Pilgrim’s Progress, Land of Beulah, …
$p_{c}$ [Uncaptioned image] Question: The book in the picture was written in the language of
$t_{c}$: English
Alias: en, eng, English language, …
@@
Question: The book in the picture is
$t^{M}_{g}$: The Pilgrim’s Progress
Alias: Pilgrim’s Progress, Land of Beulah, …
@@
Question: Which book is shown in the picture?
$t^{T}_{g}$: The Pilgrim’s Progress
Alias: Pilgrim’s Progress, Land of Beulah, …
@@
input / Question: Invisible Man was written in the language of
$\tilde{o}$: Sanskrit
$p_{r}$ / Question: Invisible Man was written in the language of
$t_{r}$: Sanskrit
Alias: Samskrta, Samskrtam, Sanskrit language, …
$p_{c}$ [Uncaptioned image] Question: The book in the picture was written in the language of
$t_{c}$: Sanskrit
Alias: Samskrta, Samskrtam, Sanskrit language, …
@@
$p^{T}_{g}$ / Question: Which language was Invisible Man written in?
$t^{T}_{g}$: Sanskrit
Alias: Samskrta, Samskrtam, Sanskrit language, …
@@
input [Uncaptioned image] Question: The official work language of the book in the picture has changed.
The book in the picture was written in the language of
$\tilde{o}$: Sanskrit
$p_{r}$ [Uncaptioned image] Question: The book in the picture was written in the language of
$t_{r}$: Sanskrit
Alias: Samskrta, Samskrtam, Sanskrit language, …
$p_{c}$ / Question: Invisible Man was written in the language of
$t_{c}$: Sanskrit
Alias: Samskrta, Samskrtam, Sanskrit language, …
@@
Question: Who is the developer of the operating system in the picture?
$t_{l}$: Microsoft
Alias: MSFT, Microsoft Corp., …
@@
Question: The book in the picture was written in the language of
$t^{M}_{g}$: Sanskrit
Alias: Samskrta, Samskrtam, Sanskrit language, …
@@
Question: Which language was the book in the picture written in?
$t^{T}_{g}$: Sanskrit
Alias: Samskrta, Samskrtam, Sanskrit language, …